Ali’s budget line relating good X and good Y has intercept of 50 unit of good X and 20 units of good Y. if the price of good X is 12, what is Ali’s income? What is the price of good Y? What is then slope of budget line?


1
Expert's answer
2021-05-26T10:30:38-0400
"I=p_x\times Q_x=12\times50=600"

"p_y=\frac {600}{20}=30"

"\frac{p_x}{p_y}=\frac{12}{30}=0.4"